Find the area of the following circles, given that:
 Diameter = 49 m

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given, diameter of circle (d) = 49 m
We know that, radius (r) = d/2
= 49/2
= 24.5 m
Then,
Area of the circle =πr^2
= 22/7 × (24.5)^2
= 22/7 × 600.25
= 22 × 85.75
= 1886.5 m^2

Thanks